OpenStudy (anonymous):

Given that x is a normally distributed random variable with a mean of 28 and a standard deviation of 7, find the following probability: P(x < 28)

OpenStudy (kropot72):

Can you find the z-score needed to solve this?

OpenStudy (anonymous):

I forgot how I keep missing up

OpenStudy (kropot72):

\[z=\frac{(x-\mu)}{\sigma}\] Just put in the values\[x=28\] \[\mu=28\]\[\sigma=7\]

OpenStudy (anonymous):

I have this z=x-μ/o P(X<28) Z=(4-28)/7=2.86 Probability is .9979 z=x-μ/o (28-28)/7=0 but my teacher said it's half right and said to find the area a to the left of this z-score.

OpenStudy (kropot72):

The correct z-score is your calculation (28 - 28)/7 = 0. When you look up the standard normal distribution table what probability do you get for z = 0.0 ?

OpenStudy (anonymous):

0.5000

OpenStudy (kropot72):

Good work! P(x < 28) = 0.500

OpenStudy (anonymous):

great thanks !!!